Exercise 6-22 (Algo) Long-term contract; revenue recognition upon project completion; loss projected on entire project [LO6-8, 6-9]

On February 1, 2021, Arrow Construction Company entered into a three-year construction contract to build a bridge for a price of $8,125,000. During 2021, costs of $2,050,000 were incurred, with estimated costs of $4,050,000 yet to be incurred. Billings of $2,560,000 were sent, and cash collected was $2,300,000.

In 2022, costs incurred were $2,560,000 with remaining costs estimated to be $3,675,000. 2022 billings were $2,810,000, and $2,525,000 cash was collected. The project was completed in 2023 after additional costs of $3,850,000 were incurred. The company’s fiscal year-end is December 31. This project does not qualify for revenue recognition over time.

Required:
1. Calculate the amount of revenue and gross profit or loss to be recognized in each of the three years.
2a. Prepare journal entries for 2021 to record the transactions described (credit "various accounts" for construction costs incurred).
2b. Prepare journal entries for 2022 to record the transactions described (credit "various accounts" for construction costs incurred).
3a. Prepare a partial balance sheet to show the presentation of the project as of December 31, 2021.
3b. Prepare a partial balance sheet to show the presentation of the project as of December 31, 2022
Answer #1

Part 1

Year

Revenue recognized

Gross profit (loss) recognized

2021

0

0

2022

0

(160000)

2023

8125000

(175000)

Total

8125000

$(335000)

Total loss = Total cost – total revenue = (2050000+2560000+3850000)-8125000 = 335000

Loss 2023 = additional costs 2023 - remaining costs estimated 2022 = 3850000-3675000 = 175000

Loss 2022 = total loss – loss 2023 = 335000-175000 = 160000

Part 2-a

No

Year

General journal

Debit

Credit

1

2021

Construction in progress

2050000

Various accounts

2050000

2

2021

Accounts receivable

2560000

Billings

2560000

3

2021

Cash

2300000

Accounts receivable

2300000

4

2021

No journal entry required

Part 2-b

No

Year

General journal

Debit

Credit

1

2022

Construction in progress

2560000

Various accounts

2560000

2

2022

Accounts receivable

2810000

Billings

2810000

3

2022

Cash

2525000

Accounts receivable

2525000

4

2022

Loss on long-term contract

160000

Construction in progress

160000

Part 3-a

Balance Sheet (Partial)

At December 31, 2021

Current assets:

Accounts receivable

260000

Current liabilities:

Billings in excess of costs

510000

Part 3-b

Balance Sheet (Partial)

At December 31, 2022

Current assets:

Accounts receivable

545000

Current liabilities:

Billings in excess of costs less loss (2560000+2810000)-(2050000+2560000-(160000))

920000